OpenStudy (anonymous):

Which improper fraction is equivalent to two and three over nine? fourteen over nine twenty-nine over nine twenty-one over nine fifteen over nine

OpenStudy (anonymous):

2 3/9 = (2 x 9 + 3)/9 = 21/9

OpenStudy (anonymous):

explain

OpenStudy (ybarrap):

$$ 2 \cfrac{3}{9} $$ means $$ 2+\cfrac{3}{9}=\cfrac{2\times9}{9}+\cfrac{3}{9}=\cfrac{2\times 9+3}{9} $$

OpenStudy (ybarrap):

make sense?

OpenStudy (anonymous):

yes